# Sequential Transport for Causal Mediation Analysis

### 16 Mar 2026 v1

@@ 1 Introduction
We integrate these transport-based ideas into causal mediation analysis. [^1] The replication package, including all data, code, and a replication ebook containing additional results, is available at: https://github.com/fer-agathe/causal_OT. For this, we consider the simple but widely used causal ordering $A\to\boldsymbol{X}\to Y,$ where $A$ is a binary treatment, $\boldsymbol{X}$ a (possibly high-dimensional) vector of mediators structured by a directed acyclic graph (DAG), and $Y$ an outcome. Our key idea is to define mediator counterfactuals in a distributional mutatis mutandis sense: when changing the treatment, each mediator is modified only as necessary to reach the target distribution, while respecting the conditional structure of the joint probability distribution encoded in the DAG.
@@ 2 Causal mediation analysis
We use uppercase letters for random variables (e.g., $A,\boldsymbol{X},Y$) and lowercase letters for their realizations (e.g., $a,\boldsymbol{x},y$); boldface denotes vectors, and $i$ indexes units/individuals while $j$ indexes mediators. Let $(A,\boldsymbol{X},Y)\in\mathcal{A}\times\mathcal{X}\times\mathcal{Y}$ be a random vector, where $A$ is a binary treatment variable with $\mathcal{A}=\{0,1\}$, $\boldsymbol{X}=(X_{1},\dots,X_{d})$ is a vector of post-treatment mediators and $Y$ is the outcome variable. We assume $(A,\boldsymbol{X},Y)\sim\mathbb{P}$, where $\mathbb{P}$ is a probability measure on the product space $\mathcal{A}\times\mathcal{X}\times\mathcal{Y}$. For each unit $i$, let $(A_{i},\boldsymbol{X}_{i},Y_{i})\sim\mathbb{P}$ denote the $i$-th random observation, and let $(a_{i},\boldsymbol{x}_{i},y_{i})$ denote its realization. We consider the causal ordering
@@
consistent with the recursive structure of a DAG, from Markov’s property (Koller and Friedman, 2009). Throughout, we distinguish the definition of sequential transport (ST) from its causal interpretation. ST is defined from the observable conditional laws $\mathbb{P}(X_{j}\mid\mathrm{parents}_{\mathcal{G}}(X_{j}),A=a)$ (or $\mathbb{P}(\cdot\mid A=a,C=c)$ if pre-treatment covariates $C$ are used) and from the mediator DAG $\mathcal{G}$; hence it yields a well-defined distributional transformation regardless of whether $A$ is exogenous. A causal interpretation requires an additional assumption: either $A$ is randomized, or $A$ is ignorable given observed pre-treatment covariates $C$, so that $\mathbb{P}(\cdot\mid A=a,C=c)=\mathbb{P}(\cdot\mid do(A=a),C=c)$. Under this condition, ST-based decompositions can be interpreted as (mutatis mutandis) direct and indirect causal effects.
To identify direct and indirect effects, we require the sequential ignorability assumption from Imai et al. (2010b), extended to multiple mediators. This entails the absence of unobserved confounding in the treatment-outcome, treatment-mediator, and mediator-outcome relationships, as well as the absence of treatment-induced confounding (see Zhou and Wodtke (2025)). When the treatment is randomized, this guarantees no confounding in the treatment-outcome and treatment-mediator relationships, but not necessarily in the other assumptions. In observational settings where $A$ is a sensitive attribute (e.g., race), it is common to treat $A$ as a root node by convention (i.e., we do not model its upstream causes). In this case, identifying causal effects additionally requires assuming that there are no unobserved variables affecting both the outcome and the mediators. When the sensitive attribute is not a root node (e.g., credit score), we must additionally control for confounding in pre-treatment variables. Thus, unless these assumptions can be justified, we do not claim identification of a causal effect of $A$ when using transport plans, in contrast to approaches based on SCMs. Rather, ST should be interpreted as a distributional pathway decomposition of differences between the $A=0$ and $A=1$ groups, aligned with $\mathcal{G}$ (see also Section 4 for a more detailed discussion).
@@ 2.1 Potential outcomes and effect decomposition
Throughout the paper, we distinguish between the definition of transport-based counterfactuals and their causal interpretation. The ST construction is purely distributional: it is built from the observed conditional laws $\mathbb{P}(X_{j}\mid\mathrm{parents}_{\mathcal{G}}(X_{j}),A=a)$ and from the structure of the mediator DAG. Hence, ST defines a transported mediator vector as soon as these conditional distributions are well-defined/estimable, regardless of whether $A$ is exogenous. A causal interpretation requires additional assumptions. First, when $A$ is a root node in the causal graph (e.g., randomized or ignorable given observed covariates), there is no unobserved confounding between treatment and outcome or between treatment and mediators. In this case, observational and interventional distributions coincide if additional sequential ignorability assumptions from Imai et al. (2010b); Zhou and Wodtke (2025), such as the ignorability of the mediators, are satisfied. Under these conditions, $\mathbb{P}(\cdot\mid A=a)=\mathbb{P}(\cdot\mid do(A=a))$ (or conditionally on covariates). In that case, the resulting ST-based decompositions admit a causal interpretation as mutatis mutandis direct and indirect effects. In particular, when $A$ is a sensitive attribute such as race in our application, we may represent $A$ as parentless in the graph by convention (that is, we do not model its upstream causes). This is a notational/expositional choice, not an exogeneity assumption. Accordingly, without further assumptions, we do not interpret the resulting decomposition as an identified causal effect of $A$; rather, it should be read as a distributional pathway decomposition of between-group differences. Moreover, in observational settings where $A$ may be confounded, ST does not claim to identify causal effects (i.e., it does not in general recover quantities under $do(A=a)$); instead, it provides a principled distributional decomposition of observed differences between the $A=0$ and $A=1$ groups aligned with the assumed mediator structure.
@@ 3.1 Transport-based counterfactuals in the literature
Hereafter, $\boldsymbol{X}=(X_{1},\dots,X_{d})$ denotes the (random) mediator vector, possibly affected by $A$ (directly or through other mediators) and influencing the outcome of interest $Y$.
@@ 3.4 Transport-induced mutatis mutandis estimands
The mapping $T_{0\to 1}$ and the estimands above are defined purely from observable distributions and thus yield a distributional decomposition of between-group differences. When $A$ is randomized (or ignorable given pre-treatment covariates), observational and interventional distributions coincide, and the same decomposition admits a causal interpretation as mutatis mutandis direct and indirect effects.
@@ 3.5 Strengths and limitations of transport-based approaches
At the same time, any transport plan from OT theory constructs counterfactuals that are distributionally identified: they are defined from the observed group distributions and can be viewed as a guided interpolation between them. In the setting considered in this paper, $A$ is assumed exogenous (a root node), so that conditioning on $A$ coincides with intervention, and the resulting transport-based decompositions admit a causal interpretation. It should be noted that in observational settings where $A$ may be confounded, transport-based counterfactuals remain well-defined as distributional objects, but interpreting them causally would require additional assumptions (e.g., ignorability given covariates) or appropriate adjustment.
@@ 4.1 Mediator DAG and notation
Let $A\in\{0,1\}$ denote the treatment and $\boldsymbol{X}=(X_{1},\dots,X_{d})$ the mediator vector. Let $\mathcal{G}$ be a mediator DAG whose nodes are $\{X_{1},\dots,X_{d}\}$ and whose directed edges encode dependencies between mediators downstream of $A$. For each mediator $X_{j}$, write $\boldsymbol{Z}_{j}=\mathrm{parents}_{\mathcal{G}}(X_{j})$ for its parent set. Because $\mathcal{G}$ is acyclic, there exists a topological ordering $X_{\pi(1)},\dots,X_{\pi(d)}$ such that all parents of $X_{\pi(j)}$ appear among $\{X_{\pi(1)},\dots,X_{\pi(j-1)}\}$ (see Kahn (1962) and for practical implementation, several algorithms are discussed in Section 20.4 in Cormen et al. (2022)).
@@ 5 Theoretical guarantees for sequential transport
Throughout, we work under assumptions compatible with the exogeneity of the treatment $A$ (a root node in the DAG) and with the absolute continuity or discrete-support structure of the mediators discussed previously. All proofs are deferred to Appendix B.
@@ 6.1 Gaussian mediation with a single mediator
where $(\varepsilon_{X},\varepsilon_{Y})$ are jointly Gaussian with zero mean and covariance matrix independent of $A$. Under this model, the natural direct and indirect effects coincide with the algebraic SEM decomposition:
@@ 6.3 Monte Carlo simulations with three mediators
and direct effects of $A$ on $Y$. In each replication, we generate $n_{0}=400$ observations with $A=0$ and $n_{1}=200$ observations with $A=1$.
@@ Results.
Figure 2 displays the empirical distributions of $\bar{\delta}$, $\bar{\zeta}$, and $\bar{\tau}$ across the $200$ Monte Carlo replications. In this mixed-Gaussian simulation setting, all transport-based approaches yield broadly similar distributions. Sequential transport (ST) shows a modest reduction in dispersion compared to multivariate optimal transport (OT) and entropic optimal transport (SKH), an effect that is more clearly seen in numerical summaries (e.g., SD/IQR) than from the violin plots alone.
@@ 7.1 Data and causal graph
We consider the standard pretrial subset originally analyzed in Kusner et al. (2017), provided with the R package fairadapt (Plečko et al., 2024), following the preprocessing conventions used in recent transport-based fairness studies (Charpentier et al., 2023; Fernandes Machado et al., 2025a, b). The outcome $Y$ is the two-year general recidivism indicator, which takes the value $1$ if a criminal defendant re-offended within two years after parole start, and 0 otherwise. The sensitive attribute is
@@
where $a=1$ denotes white defendants and $a=0$ non-white defendants. In this application, we treat $A$ as a root node by convention (we do not model its upstream causes), but we do not interpret the resulting decomposition as an identified causal effect of race; rather, it is a distributional pathway decomposition aligned with the mediator DAG.
@@ Appendix B Proofs of theoretical results
By Assumption (A3), $(I.a)\leq\displaystyle\sup_{u,\boldsymbol{z}}|\widehat{Q}_{1,j}(u\mid\boldsymbol{z})-Q_{1,j}(u\mid\boldsymbol{z})|=o_{p}(1)$. Also $\widehat{u}_{0}\xrightarrow{p}u_{0}$ by Assumption (A3); moreover, since $u_{0}\in[\varepsilon,1-\varepsilon]$, we have $\widehat{u}_{0}\in[\varepsilon/2,1-\varepsilon/2]$ with probability tending to $1$. By continuity of $u\mapsto Q_{1,j}(u\mid\widehat{\boldsymbol{z}}_{1})$ on $(0,1)$ (Assumptions (A1)(A2)), this implies $(I.b)=o_{p}(1)$. Hence $(I)=o_{p}(1)$.
@@
On the event $E_{i}$, both $\boldsymbol{x}^{\dagger}_{1,i}$ and, with probability tending to $1$, $\boldsymbol{x}^{\star}_{1,i}$ belong to $\mathcal{X}_{0}$ (since $\boldsymbol{x}^{\star}_{1,i}\to\boldsymbol{x}^{\dagger}_{1,i}$ and $\mathcal{X}_{0}$ is a support set). Then
@@ D.1 Toy example with 2 mediators
Consider the following simple example: treatment $A$ is a binary variable, taking values in $\{0,1\}$ (untreated, treated) with probabilities $p_{0}$ and $p_{1}$. Then, conditional on $A$, $\boldsymbol{X}=(X_{1},X_{2})$ is a bivariate Gaussian vector, $\boldsymbol{X}\mid A=a\sim\mathcal{N}(\boldsymbol{\mu}_{a},\boldsymbol{\Sigma}_{a})$

### 22 Mar 2026 v2

@@ 1 Introduction
We integrate these transport-based ideas into causal mediation analysis. [^1] The replication package, including all data, code, and a replication ebook containing additional results, is available at: https://github.com/fer-agathe/causal_OT. For this, we consider the simple but widely used causal ordering $A\to\boldsymbol{X}\to Y,$ where AA is a binary treatment, $\boldsymbol{X}$ a (possibly high-dimensional) vector of mediators structured by a directed acyclic graph (DAG), and $Y$ an outcome. Our key idea is to define mediator counterfactuals in a distributional mutatis mutandis sense: when changing the treatment, each mediator is modified only as necessary to reach the target distribution, while respecting the conditional structure of the joint probability distribution encoded in the DAG.
@@ 2 Causal mediation analysis
We use uppercase letters for random variables (e.g., $A,\boldsymbol{X},Y$) and lowercase letters for their realizations (e.g., $a,\boldsymbol{x},y$); boldface denotes vectors, and $i$ indexes units/individuals while $j$ indexes mediators. Let $(A,\boldsymbol{X},Y)\in\mathcal{A}\times\mathcal{X}\times\mathcal{Y}$ be a random vector, where AA is a binary treatment variable with $\mathcal{A}=\{0,1\}$, $\boldsymbol{X}=(X_{1},\dots,X_{d})$ is a vector of post-treatment mediators and $Y$ is the outcome variable. We assume $(A,\boldsymbol{X},Y)\sim\mathbb{P}$, where $\mathbb{P}$ is a probability measure on the product space $\mathcal{A}\times\mathcal{X}\times\mathcal{Y}$. For each unit $i$, let $(A_{i},\boldsymbol{X}_{i},Y_{i})\sim\mathbb{P}$ denote the $i$-th random observation, and let $(a_{i},\boldsymbol{x}_{i},y_{i})$ denote its realization. We consider the causal ordering
@@
consistent with the recursive structure of a DAG, from Markov’s property (Koller and Friedman, 2009). Throughout, we distinguish the definition of sequential transport (ST) from its causal interpretation. ST is defined from the observable conditional laws $\mathbb{P}(X_{j}\mid\mathrm{parents}_{\mathcal{G}}(X_{j}),A=a)$ (or $\mathbb{P}(\cdot\mid A=a,C=c)$ if pre-treatment covariates $C$ are used) and from the mediator DAG $\mathcal{G}$; hence it yields a well-defined distributional transformation regardless of whether AA is exogenous. A causal interpretation requires an additional assumption: either AA is randomized, or AA is ignorable given observed pre-treatment covariates $C$, so that $\mathbb{P}(\cdot\mid A=a,C=c)=\mathbb{P}(\cdot\mid do(A=a),C=c)$. Under this condition, ST-based decompositions can be interpreted as (mutatis mutandis) direct and indirect causal effects.
To identify direct and indirect effects, we require the sequential ignorability assumption from Imai et al. (2010b), extended to multiple mediators. This entails the absence of unobserved confounding in the treatment-outcome, treatment-mediator, and mediator-outcome relationships, as well as the absence of treatment-induced confounding (see Zhou and Wodtke (2025)). When the treatment is randomized, this guarantees no confounding in the treatment-outcome and treatment-mediator relationships, but not necessarily in the other assumptions. In observational settings where AA is a sensitive attribute (e.g., race), it is common to treat AA as a root node by convention (i.e., we do not model its upstream causes). In this case, identifying causal effects additionally requires assuming that there are no unobserved variables affecting both the outcome and the mediators. When the sensitive attribute is not a root node (e.g., credit score), we must additionally control for confounding in pre-treatment variables. Thus, unless these assumptions can be justified, we do not claim identification of a causal effect of AA when using transport plans, in contrast to approaches based on SCMs. Rather, ST should be interpreted as a distributional pathway decomposition of differences between the $A=0$ and $A=1$ groups, aligned with $\mathcal{G}$ (see also Section 4 for a more detailed discussion).
@@ 2.1 Potential outcomes and effect decomposition
Throughout the paper, we distinguish between the definition of transport-based counterfactuals and their causal interpretation. The ST construction is purely distributional: it is built from the observed conditional laws $\mathbb{P}(X_{j}\mid\mathrm{parents}_{\mathcal{G}}(X_{j}),A=a)$ and from the structure of the mediator DAG. Hence, ST defines a transported mediator vector as soon as these conditional distributions are well-defined/estimable, regardless of whether AA is exogenous. A causal interpretation requires additional assumptions. First, when AA is a root node in the causal graph (e.g., randomized or ignorable given observed covariates), there is no unobserved confounding between treatment and outcome or between treatment and mediators. In this case, observational and interventional distributions coincide if additional sequential ignorability assumptions from Imai et al. (2010b); Zhou and Wodtke (2025), such as the ignorability of the mediators, are satisfied. Under these conditions, $\mathbb{P}(\cdot\mid A=a)=\mathbb{P}(\cdot\mid do(A=a))$ (or conditionally on covariates). In that case, the resulting ST-based decompositions admit a causal interpretation as mutatis mutandis direct and indirect effects. In particular, when AA is a sensitive attribute such as race in our application, we may represent AA as parentless in the graph by convention (that is, we do not model its upstream causes). This is a notational/expositional choice, not an exogeneity assumption. Accordingly, without further assumptions, we do not interpret the resulting decomposition as an identified causal effect of AA; rather, it should be read as a distributional pathway decomposition of between-group differences. Moreover, in observational settings where AA may be confounded, ST does not claim to identify causal effects (i.e., it does not in general recover quantities under $do(A=a)$); instead, it provides a principled distributional decomposition of observed differences between the $A=0$ and $A=1$ groups aligned with the assumed mediator structure.
@@ 3.1 Transport-based counterfactuals in the literature
Hereafter, $\boldsymbol{X}=(X_{1},\dots,X_{d})$ denotes the (random) mediator vector, possibly affected by AA (directly or through other mediators) and influencing the outcome of interest $Y$.
@@ 3.4 Transport-induced mutatis mutandis estimands
The mapping $T_{0\to 1}$ and the estimands above are defined purely from observable distributions and thus yield a distributional decomposition of between-group differences. When AA is randomized (or ignorable given pre-treatment covariates), observational and interventional distributions coincide, and the same decomposition admits a causal interpretation as mutatis mutandis direct and indirect effects.
@@ 3.5 Strengths and limitations of transport-based approaches
At the same time, any transport plan from OT theory constructs counterfactuals that are distributionally identified: they are defined from the observed group distributions and can be viewed as a guided interpolation between them. In the setting considered in this paper, AA is assumed exogenous (a root node), so that conditioning on AA coincides with intervention, and the resulting transport-based decompositions admit a causal interpretation. It should be noted that in observational settings where AA may be confounded, transport-based counterfactuals remain well-defined as distributional objects, but interpreting them causally would require additional assumptions (e.g., ignorability given covariates) or appropriate adjustment.
@@ 4.1 Mediator DAG and notation
Let $A\in\{0,1\}$ denote the treatment and $\boldsymbol{X}=(X_{1},\dots,X_{d})$ the mediator vector. Let $\mathcal{G}$ be a mediator DAG whose nodes are $\{X_{1},\dots,X_{d}\}$ and whose directed edges encode dependencies between mediators downstream of AA. For each mediator $X_{j}$, write $\boldsymbol{Z}_{j}=\mathrm{parents}_{\mathcal{G}}(X_{j})$ for its parent set. Because $\mathcal{G}$ is acyclic, there exists a topological ordering $X_{\pi(1)},\dots,X_{\pi(d)}$ such that all parents of $X_{\pi(j)}$ appear among $\{X_{\pi(1)},\dots,X_{\pi(j-1)}\}$ (see Kahn (1962) and for practical implementation, several algorithms are discussed in Section 20.4 in Cormen et al. (2022)).
@@ 5 Theoretical guarantees for sequential transport
Throughout, we work under assumptions compatible with the exogeneity of the treatment AA (a root node in the DAG) and with the absolute continuity or discrete-support structure of the mediators discussed previously. All proofs are deferred to Appendix B.
@@ 6.1 Gaussian mediation with a single mediator
where $(\varepsilon_{X},\varepsilon_{Y})$ are jointly Gaussian with zero mean and covariance matrix independent of AA. Under this model, the natural direct and indirect effects coincide with the algebraic SEM decomposition:
@@ 6.3 Monte Carlo simulations with three mediators
and direct effects of AA on $Y$. In each replication, we generate $n_{0}=400$ observations with $A=0$ and $n_{1}=200$ observations with $A=1$.
@@ Results.
Figure 2 displays the empirical distributions of $\bar{\delta}$, $\bar{\zeta}$, and $\bar{\tau}$ across the 200200 Monte Carlo replications. In this mixed-Gaussian simulation setting, all transport-based approaches yield broadly similar distributions. Sequential transport (ST) shows a modest reduction in dispersion compared to multivariate optimal transport (OT) and entropic optimal transport (SKH), an effect that is more clearly seen in numerical summaries (e.g., SD/IQR) than from the violin plots alone.
@@ 7.1 Data and causal graph
We consider the standard pretrial subset originally analyzed in Kusner et al. (2017), provided with the R package fairadapt (Plečko et al., 2024), following the preprocessing conventions used in recent transport-based fairness studies (Charpentier et al., 2023; Fernandes Machado et al., 2025a, b). The outcome $Y$ is the two-year general recidivism indicator, which takes the value 11 if a criminal defendant re-offended within two years after parole start, and 0 otherwise. The sensitive attribute is
@@
where $a=1$ denotes white defendants and $a=0$ non-white defendants. In this application, we treat AA as a root node by convention (we do not model its upstream causes), but we do not interpret the resulting decomposition as an identified causal effect of race; rather, it is a distributional pathway decomposition aligned with the mediator DAG.
@@ Appendix B Proofs of theoretical results
By Assumption (A3), $(I.a)\leq\displaystyle\sup_{u,\boldsymbol{z}}|\widehat{Q}_{1,j}(u\mid\boldsymbol{z})-Q_{1,j}(u\mid\boldsymbol{z})|=o_{p}(1)$. Also $\widehat{u}_{0}\xrightarrow{p}u_{0}$ by Assumption (A3); moreover, since $u_{0}\in[\varepsilon,1-\varepsilon]$, we have $\widehat{u}_{0}\in[\varepsilon/2,1-\varepsilon/2]$ with probability tending to 11. By continuity of $u\mapsto Q_{1,j}(u\mid\widehat{\boldsymbol{z}}_{1})$ on $(0,1)$ (Assumptions (A1)(A2)), this implies $(I.b)=o_{p}(1)$. Hence $(I)=o_{p}(1)$.
@@
On the event $E_{i}$, both $\boldsymbol{x}^{\dagger}_{1,i}$ and, with probability tending to 11, $\boldsymbol{x}^{\star}_{1,i}$ belong to $\mathcal{X}_{0}$ (since $\boldsymbol{x}^{\star}_{1,i}\to\boldsymbol{x}^{\dagger}_{1,i}$ and $\mathcal{X}_{0}$ is a support set). Then
@@ D.1 Toy example with 2 mediators
Consider the following simple example: treatment AA is a binary variable, taking values in $\{0,1\}$ (untreated, treated) with probabilities $p_{0}$ and $p_{1}$. Then, conditional on AA, $\boldsymbol{X}=(X_{1},X_{2})$ is a bivariate Gaussian vector, $\boldsymbol{X}\mid A=a\sim\mathcal{N}(\boldsymbol{\mu}_{a},\boldsymbol{\Sigma}_{a})$